### v2

@@ 1 Introduction
The buyer of a vanilla basket option of strike $K$ and maturity $T$ on underlying assets $(S_{i})_{i=1,...,n}$ receives at maturity $T$ $\max\left(\sum_{i=1}^{n}w_{i}S_{i}(T)-K\right)$ for a call and $\max\left(K-\sum_{i=1}^{n}w_{i}S_{i}(T)\right)$ for a put. The weights $w_{i}$ typically sum to one. The buyer of an Asian call option of strike $K$ and averaging dates $(t_{i})_{i=1,...,n}$ on an asset $S$ receives $\max\left(\sum_{i=1}^{n}w_{i}S(t_{i})-K\right)$. Under the Black-Scholes model with a term-structure of interest rates, dividend yields and volatilities, Asian options with discrete arithmetic averaging can be reduced to an equivalent vanilla basket option with a specific correlation structure. Asian (basket) options are particularly popular among commodity traders, but is also traded regularly in the foreign exchange and in the equity derivatives spaces. There is no exact closed analytic formula for the pricing of arithmetic Asian or vanilla basket options. The problem has been widely studied. In the context of the Black-Scholes model, early valuation techniques were based on approximating the arithmetic mean by a geometric mean [17], or on moment matching against a lognormal distribution as in Levy [12]. Later, Curran [4] introduced a much more accurate pricing based on geometric conditioning which has been more recently further improved, by exploring better integration cutoffs and conditioning variables [5]. Ju proposed a third-order correction for the Levy moment matching in [9], while Zhou [19] looked at matching four moments with the log extended skew normal distribution. Beyond the Black-Scholes model, but restricted to Asian options, Zhang and Oosterlee propose an efficient method based on the Fourier cosine series for affine models with a known characteristic function in [18]. Here, we consider a Black-Scholes setting with time-dependent parameters, and propose a stochastic expansion using the geometric average as a proxy, following the ideas of Etoré and Gobet, who applied their expansion to price vanilla options with discrete dividends under the piecewise-lognormal model in [6], with further refinements in [11]. We start by deriving the first, second and third-order expansions and then evaluate their accuracy on a sample of previously published cases. This allows for a comparison with other accurate approximations.
@@ 2.1 From Asian to basket
$F_{i}(0,T)=\mathbb{E}[S_{i}(T)]=\mathbb{E}[S(t_{i})]=F(0,t_{i})=S(0)e^{\int_{0}^{t_{i}}(r(s)-q(s))\mathop{}\!\kern 0.0pt\mathrm{d}s}\,.$ (2)
@@ 2.2 Asian baskets
$\hat{\rho}_{i+n(j-1),k+n(l-1)}=\rho_{j,l}\frac{\int_{0}^{t_{i}\wedge t_{k}}\sigma_{j}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\int_{0}^{t_{k}}\sigma_{l}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}}\,,$ (3)
@@ 3.1 The proxy
$\tilde{a}_{i}=\frac{w_{i}}{A(T)}\,,\quad A=\sum_{i=1}^{n}w_{i}F_{i}(0,T)\,,\quad{K}^{\star}=\frac{K}{A}\,,$
$S_{i}^{\star}(t)=\frac{S(t)}{F(0,t)}=e^{-\frac{1}{2}\int_{0}^{t}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}\,.$
@@
$V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\\ \right]$ (4)
@@
$m(T)=\mathbb{E}\left[\ln G(T)\right]$ $=-\frac{1}{2}\sum_{i=1}^{n}a_{i}v_{i}\,,$ (5)
$\tilde{\nu}^{2}(T)=\mathrm{Var}\left[\ln G(T)\right]$ $=\sum_{i,j=1}^{n}a_{i}a_{j}\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$ (6)
with $v_{i}=\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$. Vorst [17], Gentle [7] propose to adjust the strike price of the geometric average option such that the expectation of the term inside the max function is preserved and use $\tilde{K}^{\star}=K^{\star}-(1-\mathbb{E}\left[G(T)\right])$ instead of $K^{\star}$. A different way to preserve the expectation consists in adjusting the spot side instead of the strike side:
$V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\right]$ (7)
@@
$\nu_{A}^{2}=\ln\left(\mathrm{Var}\left[A(T)\right]\right)=\ln\left(\sum_{i,j=1}^{n}a_{i}a_{j}e^{\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\right)\,..$ (8)
@@
$a_{i}=\tilde{a}_{i}\frac{\nu_{A}}{\tilde{\nu}}\,.$ (9)
@@
$V(0)\approx A\cdot\mathrm{Black}(1,K^{\star},{\nu}^{2},T)$ (10)
@@
$\mathrm{Black}(F,K,v,T)=\eta B(0,T)\left[F\Phi(\eta d_{1})-K\Phi(\eta d_{2})\right]$ (11)
@@
We will adopt the Vorst geometric formula with adjusted forward as proxy for our stochastic Taylor expansion. The Vorst geometric formula with strike shift leads to slightly more complicated equations. Interestingly, it results in exactly the same approximation for expansions of order 1 and above.
@@ 3.2 First order expansion
$\mathbb{E}\left[B(0,T)h\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-K^{\star}\right)\right]=\mathbb{E}\left[B(0,T)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathbb{E}\left[B(0,T)\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathrm{Error}_{2}(h)$ (12)
@@
$\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ =-\frac{\partial}{\partial K^{\star}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]$ (13)
@@
$\mathbb{E}\left[G^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{G}\left[h\left(\tilde{G}^{\star}(T)e^{\nu^{2}(T)}-K^{\star}\right)\right]\,.$ (14)
@@
$\mathbb{E}\left[S_{i}^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{i}\left[h\left(\tilde{G}^{\star}(T)e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]\,.$ (15)
@@
$V_{\textsf{VG1}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\,.$ (16)
@@
$\bar{v}_{i}=\sum_{l=1}^{n}a_{l}v_{i,l}\,,$ $\quad v_{i,l}=\rho_{i,l}\int_{0}^{T}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$ (17)
@@ 3.3 Second order expansion
$V_{\textsf{VG2}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ +\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)+\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ -A\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)\,.$ (18)
@@ 3.4 Third order expansion
$V_{\textsf{VG3}}$ $=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)$ (19) $+A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)$ $-A\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{6}e^{2\nu^{2}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{3\nu^{2}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\nu^{2}+2\bar{v}_{i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{2\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{\bar{v}_{i}+\bar{v}_{j}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}+v_{j}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{6}\sum_{i,j,l=1}^{n}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\,,$

### v3

@@ 1 Introduction
The buyer of a vanilla basket option of strike $K$ and maturity $T$ on underlying assets $(S_{i})_{i=1,...,n}$ receives at maturity $T$ $\max\left(\sum_{i=1}^{n}w_{i}S_{i}(T)-K\right)$ for a call and $\max\left(K-\sum_{i=1}^{n}w_{i}S_{i}(T)\right)$ for a put. The weights $w_{i}$ typically sum to one. The buyer of an Asian call option of strike $K$ and averaging dates $(t_{i})_{i=1,...,n}$ on an asset $S$ receives $\max\left(\sum_{i=1}^{n}w_{i}S(t_{i})-K\right)$. Under the Black-Scholes model with a term-structure of interest rates, dividend yields and volatilities, Asian options with discrete arithmetic averaging can be reduced to an equivalent vanilla basket option with a specific correlation structure. Asian (basket) options are particularly popular among commodity traders, but is also traded regularly in the foreign exchange and in the equity derivatives spaces.
There is no exact closed analytic formula for the pricing of arithmetic Asian or vanilla basket options. The problem has been widely studied. In the context of the Black-Scholes model, early valuation techniques were based on approximating the arithmetic mean by a geometric mean [17], or on moment matching against a lognormal distribution as in Levy [12]. Later, Curran [4] introduced a much more accurate pricing based on geometric conditioning which has been more recently further improved, by exploring better integration cutoffs and conditioning variables [5]. Ju proposed a third-order correction for the Levy moment matching in [9], while Zhou [19] looked at matching four moments with the log extended skew normal distribution. Beyond the Black-Scholes model, but restricted to Asian options, Zhang and Oosterlee propose an efficient method based on the Fourier cosine series for affine models with a known characteristic function in [18].
Here, we consider a Black-Scholes setting with time-dependent parameters, and propose a stochastic expansion using the geometric average as a proxy, following the ideas of Etoré and Gobet, who applied their expansion to price vanilla options with discrete dividends under the piecewise-lognormal model in [6], with further refinements in [11]. We start by deriving the first, second and third-order expansions and then evaluate their accuracy on a sample of previously published cases. This allows for a comparison with other accurate approximations.
@@ 2.1 From Asian to basket
(2) $F_{i}(0,T)=\mathbb{E}[S_{i}(T)]=\mathbb{E}[S(t_{i})]=F(0,t_{i})=S(0)e^{\int_{0}^{t_{i}}(r(s)-q(s))\mathop{}\!\kern 0.0pt\mathrm{d}s}\,.$
@@ 2.2 Asian baskets
(3) $\hat{\rho}_{i+n(j-1),k+n(l-1)}=\rho_{j,l}\frac{\int_{0}^{t_{i}\wedge t_{k}}\sigma_{j}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}{\sqrt{\int_{0}^{t_{i}}\sigma_{j}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\int_{0}^{t_{k}}\sigma_{l}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}}\,,$
@@ 3.1 The proxy
$\tilde{a}_{i}=\frac{w_{i}F_{i}(0,T)}{A(T)}\,,\quad A=\sum_{i=1}^{n}w_{i}F_{i}(0,T)\,,\quad{K}^{\star}=\frac{K}{A}\,,$
$S_{i}^{\star}(t)=\frac{S_{i}(t)}{F_{i}(0,t)}=e^{-\frac{1}{2}\int_{0}^{t}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s+\int_{0}^{t}\sigma_{i}(s)\mathop{}\!\kern 0.0pt\mathrm{d}W_{i}(s)}\,.$
@@
(4) $V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\\ \right]$
@@
(5) $m(T)=\mathbb{E}\left[\ln G(T)\right]$ $=-\frac{1}{2}\sum_{i=1}^{n}a_{i}v_{i}\,,$
(6) $\tilde{\nu}^{2}(T)=\mathrm{Var}\left[\ln G(T)\right]$ $=\sum_{i,j=1}^{n}a_{i}a_{j}\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
with $v_{i}=\int_{0}^{T}\sigma_{i}^{2}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s$.
Vorst [17], Gentle [7] propose to adjust the strike price of the geometric average option such that the expectation of the term inside the max function is preserved and use $\tilde{K}^{\star}=K^{\star}-(1-\mathbb{E}\left[G(T)\right])$ instead of $K^{\star}$. A different way to preserve the expectation consists in adjusting the spot side instead of the strike side:
(7) $V(0)\approx B(0,T)A\mathbb{E}\left[\left(\eta\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-{K}^{\star}\right)\right)^{+}\right]$
@@
(8) $\nu_{A}^{2}=\ln\left(\mathrm{Var}\left[A(T)\right]\right)=\ln\left(\sum_{i,j=1}^{n}a_{i}a_{j}e^{\rho_{i,j}\int_{0}^{T}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}\right)\,..$
@@
(9) $a_{i}=\tilde{a}_{i}\frac{\nu_{A}}{\tilde{\nu}}\,.$
@@
(10) $V(0)\approx A\cdot\mathrm{Black}(1,K^{\star},{\nu}^{2},T)$
@@
(11) $\mathrm{Black}(F,K,v,T)=\eta B(0,T)\left[F\Phi(\eta d_{1})-K\Phi(\eta d_{2})\right]$
@@
We will adopt the Vorst geometric formula with adjusted forward as proxy for our stochastic Taylor expansion. The Vorst geometric formula with strike shift would lead to slightly more complicated equations and exhibit a worse accuracy.
@@ 3.2 First order expansion
(12) $\mathbb{E}\left[B(0,T)h\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-K^{\star}\right)\right]=\mathbb{E}\left[B(0,T)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathbb{E}\left[B(0,T)\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ +\mathrm{Error}_{2}(h)$
@@
(13) $\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h^{\prime}\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]\\ =-\frac{\partial}{\partial K^{\star}}\mathbb{E}\left[\left(\sum_{i=1}^{n}\tilde{a}_{i}S_{i}^{\star}(T)-\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}\right)h\left(\alpha\prod_{i=1}^{n}S_{i}^{\star}(T)^{a_{i}}-K^{\star}\right)\right]$
@@
(14) $\mathbb{E}\left[G^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{G}\left[h\left(\tilde{G}^{\star}(T)e^{\nu^{2}(T)}-K^{\star}\right)\right]\,.$
@@
(15) $\mathbb{E}\left[S_{i}^{\star}(T)h^{\prime}\left(G^{\star}(T)-K^{\star}\right)\right]$ $=-\frac{\partial}{\partial K^{\star}}\mathbb{E}^{i}\left[h\left(\tilde{G}^{\star}(T)e^{\sum_{j=1}^{n}a_{j}\int_{0}^{T}\rho_{i,j}\sigma_{i}(s)\sigma_{j}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s}-K^{\star}\right)\right]\,.$
@@
(16) $V_{\textsf{VG1}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\,.$
@@
(17) $\bar{v}_{i}=\sum_{l=1}^{n}a_{l}v_{i,l}\,,$ $\quad v_{i,l}=\rho_{i,l}\int_{0}^{T}\sigma_{i}(s)\sigma_{l}(s)\mathop{}\!\kern 0.0pt\mathrm{d}s\,.$
@@ 3.3 Second order expansion
(18) $V_{\textsf{VG2}}=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)\\ +A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ +\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)+\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)\\ -A\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)\,.$
@@ 3.4 Third order expansion
(19) $V_{\textsf{VG3}}$ $=A\cdot\mathrm{Black}(1,K^{\star},\nu^{2},T)$ $+A\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\nu^{2}},K^{\star},\nu^{2},T\right)-A\sum_{i=1}^{n}\tilde{a}_{i}\frac{\partial}{\partial K^{\star}}\mathrm{Black}\left(e^{\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}e^{\nu^{2}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},K^{\star},\nu^{2},T\right)$ $-A\sum_{i=1}^{n}\tilde{a}_{i}e^{\bar{v}_{i}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{v_{i,j}}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{6}e^{2\nu^{2}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{3\nu^{2}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{2}\sum_{i=1}^{n}\tilde{a}_{i}e^{\nu^{2}+2\bar{v}_{i}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{2\nu^{2}+\bar{v}_{i}},K^{\star},\nu^{2},T\right)$ $+\frac{A}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}\tilde{a}_{j}e^{\bar{v}_{i}+\bar{v}_{j}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\nu^{2}+\bar{v}_{i}+v_{j}},K^{\star},\nu^{2},T\right)$ $-\frac{A}{6}\sum_{i,j,l=1}^{n}\tilde{a}_{i}\tilde{a}_{j}\tilde{a}_{l}e^{v_{i,l}+v_{j,l}+v_{i,j}}\frac{\partial^{3}}{\partial{K^{\star}}^{3}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{l}},K^{\star},\nu^{2},T\right)\,,$